10-3 Hillside issues $2,100,000 of 5%, 15-year bonds dated January 1, 2017, that pay interest semiannually...

10-3

Hillside issues $2,100,000 of 5%, 15-year bonds dated January 1, 2017, that pay interest semiannually on June 30 and December 31. The bonds are issued at a price of $2,570,390.


Required:

1. Prepare the January 1, 2017, journal entry to record the bonds’ issuance.
2(a) For each semiannual period, complete the table below to calculate the cash payment.
2(b) For each semiannual period, complete the table below to calculate the straight-line premium amortization.
2(c) For each semiannual period, complete the table below to calculate the bond interest expense.
3. Complete the below table to calculate the total bond interest expense to be recognized over the bonds' life.
4. Prepare the first two years of an amortization table using the straight-line method
5. Prepare the journal entries to record the first two interest payments.

For each semiannual period, complete the table below to calculate the cash payment, straight-line premium amortization and bond interest expense. (Round "Unamortized Premium" to whole dollar and use the rounded value for part 4 & 5.)

Par (maturity) value Annual Rate Year Semiannual cash interest payment
not attempted not attempted not attempted not attempted not attempted = not attempted
Bond price Par (maturity value) Premium on Bonds Payable Semiannual periods Straight-line premium amortization
not attempted not attempted not attempted = not attempted not attempted not attempted = not attempted
Semiannual cash payment Premium amortization Bond interest expense
not attempted not attempted not attempted = not attempted
Answer #1
Journal entry
Date General Journal Debit Credit
1/1/2018 Cash 2,570,390
premium on bonds 470,390
bonds payable 2,100,000
2-a) par maturity value Annual rate / year semi annual cash payment
2,100,000 * 5% 6./12 52500
semi annual Straight line
2-b) bond price par value premium periods premium amortization
2,570,390 - 2,100,000 = 470,390 / 30 = 15680
2-c) Semi annual cash premium bond interest expense
payment amortization
52,500 - 15680 = 36,820
3) total bond interest expense over life of bonds
amount repaid
30 payments of 52,500 1575000
par value ant maturity 2,100,000
total repaid 3675000
less amount borrowed 2,570,390
total bond interest expense. 1,104,610
(note bond interest expense may differ slightly due to rounding)
4) unamort Carrying
period premium value
1/1/2018 470,390 2,570,390
06/30/18 454,710 2,554,710
12/31/18 439,031 2,539,031
06/30/19 423,351 2,523,351
12/31/19 407,671 2,507,671
5)
Date General Journal Debit Credit
6/30/18 interest expense 36,820
premium on bonds payable 15,680
cash 52,500
31/12/2018
interest expense 36,820
premium on bonds payable 15,680
cash 52,500